. The oxygen: air entrainment ratio for 60% oxygen is ___________

2. Calculate the air: O2 ratio of 30% O2: air entrainment device ___________

3. An oxygen flowrate of 12 L/min is used on a 60% Oxygen: air entrainment mask. What is the O2: air entrainment ratio and total flow? ___________

4. If the patient's blanket is covering the port on the air entrainment device, what is the FiO2 that is being delivered? _______%

5. Which of these methods cause sterilization?

7. You want to achieve an oxygen concentration of 80% by blending air and O2 together. To achieve this, you would have an air:O2 ratio of _______.

8. The Henderson-Hasselbalch equation states that pH = ___________.

9. Four pyrrole rings form a _____________.

10. A left-shifted oxyhemoglobin dissociation curve is associated with _______________.

11. The PO2 necessary to obtain 50% hemoglobin saturation is _______________.

12. If hemoglobin saturation is kept above _______%, the PO2 can change a lot with minimal change in saturation.

13. Based on morphology as viewed through a light microscope, bacteria that are cylindrical or rod-shaped are called ______________.

14. In the Gram stain procedure, bacteria that remain blue after the counterstain has been applied are called ___________________.

15. Which statements refer to normal microbial flora?

16. What structures are responsible for bacterial movement? ______________

17. What is a nosocomial infection? _________________

18. Which type of bacteria are more likely to be MRSA? _________________

19. Because mycobacteria's cell walls retain carbolfuchsin stain despite rinsing with hydrochloric acid, they are called _______________.

20. Identify two differences of viruses and bacteria.
Answer #1

Answers are:

1. Air/oxygen ratio- 100- O2%/ O2%- 21   ( oxygen in air=21%)   

- 100- 60/ 60-21

- 40/ 39 = 1:1

2. Using above formula- 100- 30/30-21

- 70/9 = 7.8 =8:1

3. Using above formula- 100-60/60-21 = 40/39= 1:1

Total flow calculation= 1:1 =1+1 =2 (multiply with Litres per minute ie. 12)

So, 2X 12= 36

7. Using air/ oxygen ratio formula= 100-80/ 80-21

= 20/ 59 = 0.3:1